Find the thickness of the material:

This latest alien sample is crazy thin and very tough. After it was cut into something very rectangle-ish

with sides of 54.2 cm and 12.3 cm. The calculated density of the material is 1.3 g/cm^3. The mass was

found to be 11.4 grams.

Find the thickness of the alien material,

(if the answer is less than 1, then make sure you put a 'o' in front of the decimal).

(Round to hundredths place).

The thickness or the height of the material is

om 


1
Expert's answer
2021-03-04T06:42:54-0500

Volume=Mass/Density=11.4/1.3=8.7692cm3

V=l*w*h

8.7692=54.2*12.3*h

h=8.7692/(54.2*12.3)

h=0.01315cm

=0.01cm